# An Analytic Computation of Three–Loop Five–Point Feynman Integrals

### v1

@@ I Introduction
One of the challenging parts of a multi–loop perturbative computation is the calculation of the appearing Feynman integrals. Significant progress has been made in this direction, with innovative techniques developed, such as differential equations Henn:2013pwa , sector decomposition Borowka:2017idc , auxiliary mass flow Liu:2022chg , difference equation Lee:2012te , etc.. In particular, many multi–loop Feynman integrals with four, five, and even six external legs, important for phenomenology and theory, were calculated analytically Chicherin:2018old ; Abreu:2018rcw ; Chicherin:2020oor ; Abreu:2020jxa ; Canko:2020ylt ; Abreu:2021smk ; Chicherin:2021dyp ; Abreu:2023rco ; Abreu:2024yit ; Jiang:2024eaj ; Henn:2023vbd ; DiVita:2014pza ; Long:2024bmi ; Henn:2021cyv ; Henn:2024ngj . The analytic expressions, when coded in public packages, make the phenomenology computation highly efficient. Analytic results of Feynman integrals also provide the basis for theoretical studies in perturbative quantum field theory.
@@
In order to get the weight–six solution to the differential equation, we introduce a novel iterated integral representation. The solutions at weights four to six are all expressed as one–fold integrations over products of classical polylogarithms. Our solutions agree with standard numerical program pySecDec Borowka:2017idc ; Li:2015foa ; Borowka:2018goh for all integrals, and with AMFlow Liu:2022chg for the integrals calculable in the current version of AMFlow.
@@ VIII Auxiliary Matrix for Higher Weight integration
With the analytic weight three solutions and the boundary constants at hand, we can turn towards the computation of the integrals up to weight six. However, beyond weight three, classical polylogarithms are no longer sufficient and the number and complexity of the functions grow significantly. For multi–scale problems with square roots, obtaining higher weight expressions in terms of MPLs is still a highly non–trivial task. Moreover, such explicit MPL representations are not always good for efficient numerical evaluations. Therefore, it is more practical to express the higher weight iterative integrals in terms of a one–fold integral representation Caron-Huot:2014lda ; Gehrmann:2018yef ; Chicherin:2021dyp ; Henn:2024ngj .

### v2

@@ I Introduction
One of the challenging parts of a multi–loop perturbative computation is the calculation of the appearing Feynman integrals. Significant progress has been made in this direction, with innovative techniques developed, such as differential equations Henn:2013pwa , sector decomposition Borowka:2017idc , auxiliary mass flow Liu:2022chg ; Liu:2017jxz , difference equation Lee:2012te , etc.. In particular, many multi–loop Feynman integrals with four, five, and even six external legs, important for phenomenology and theory, were calculated analytically Chicherin:2018old ; Abreu:2018rcw ; Chicherin:2020oor ; Abreu:2020jxa ; Canko:2020ylt ; Abreu:2021smk ; Chicherin:2021dyp ; Abreu:2023rco ; Abreu:2024yit ; Jiang:2024eaj ; Henn:2023vbd ; DiVita:2014pza ; Long:2024bmi ; Henn:2021cyv ; Henn:2024ngj ; Chen:2022yni ; Chen:2022pdw . The analytic expressions, when coded in public packages, make the phenomenology computation highly efficient. Analytic results of Feynman integrals also provide the basis for theoretical studies in perturbative quantum field theory.
@@
In order to get the weight–six solution to the differential equation, we introduce a novel iterated integral representation. The solutions at weights four to six are all expressed as one–fold integrations over products of classical polylogarithms. Our solutions agree with standard numerical program pySecDec Borowka:2017idc ; Li:2015foa ; Borowka:2018goh for all integrals, and with AMFlow Liu:2022chg ; Liu:2017jxz for the integrals calculable in the current version of AMFlow.
@@ VIII Auxiliary Matrix for Higher Weight integration
With the analytic weight three solutions and the boundary constants at hand, we can turn towards the computation of the integrals up to weight six. However, beyond weight three, classical polylogarithms are no longer sufficient and the number and complexity of the functions grow significantly. For multi–scale problems with square roots, obtaining higher weight expressions in terms of MPLs is still a highly non–trivial task. Moreover, such explicit MPL representations are not always good for efficient numerical evaluations. Therefore, it is more practical to express the higher weight iterative integrals in terms of a one–fold integral representation Caron-Huot:2014lda ; Gehrmann:2018yef ; Chicherin:2021dyp ; Wang:2022enl ; Henn:2024ngj .